OpenStudy (anonymous):

solve 6(3 + 4x ) - 5 > greater than or equal to 181 the solution set is { x | x >greater than or equal to what answer }

OpenStudy (anonymous):

\[6(3+4x) - 5 \ge 181\]\[18 + 24x - 5 \ge 181\]\[13 + 24x \ge 181\]\[24x \ge168\]\[x \ge 7 \]

OpenStudy (anonymous):

thanks again rainbullet you're too smart